# Workload Estimation for Unknown Tasks: A Survey of Machine Learning Under Distribution Shift

### v1

@@ III Notation and Evaluation Criteria
Portability refers to the number of distributions used to train and test a model [56]. Learning tasks may consist of multiple source domains (e.g., rural hospital data vs. urban hospital data). A straightforward means of combining information from multiple domains is to construct a single joint distribution; however, data in some domains may be of higher quality than others. Weighting high-quality data over low-quality data is appropriate to maximize performance [56]. Additionally, the separation of source domains may be artificial. Prior work established that modeling domains separately can be beneficial for learning inductive biases that improve performance on unknown learning tasks [57]. Portability can have four values: one-to-one, one-to-many, many-to-one, and many-to-many. One-to-one refers to an algorithm that transfers knowledge between one source and one target domain. One-to-many transfers knowledge from one source domain to many target domains, and many-to-one transfers knowledge from multiple source domains to a single target domain. Many-to-many knowledge transfer occurs between multiple sources to multiple target domains simultaneously.
@@
Figure 1: A high level overview of non-IID machine learning methods. Adapted from Figure 19.11 of [2].
@@
Model Complexity refers to the number of parameters an underlying machine learning model needs for the algorithm to be successful [58]. Machine learning models with more parameters require higher volumes of training data. Early statisticians built simple linear models using hundreds of examples [58], whereas modern deep learning architectures may require tens of billions of examples [59]. Model complexity shares a relationship with both the underlying machine learning model type and problem complexity, as more complex models are often required to solve more complex problems. Workload estimation and many HRI problems are typically conducted in low-data regimes, as they require human subject evaluations for machine learning model development; thus, smaller models are preferred. An example of low model complexity is a random forest model that contains a few hundred decision trees, where the depth of each tree is dataset-dependent; thus, these models typically contain on the order of thousands of parameters [60]. An example of medium model complexity is shallow neural networks (i.e., fully connected neural networks) that typically contain hundreds of thousands of parameters [22], but may contain on the order of millions of parameters. An example of high model complexity is deep neural networks (e.g., convolutional networks, long short-term memory networks (LSTM)), also referred to as deep learning architectures. Deep learning architectures can contain anywhere from tens of millions [61] to over 100 billion parameters [59].
@@
The last metric, Adaptability, refers to the amount of data required to update a machine learning model to solve new problems [62] [63]. Many non-IID machine learning methods seek to account for distribution shifts by updating existing machine learning models’ parameters so that the model can solve new learning tasks. Adaptability is the number of test data points necessary for the updated model to make accurate predictions for new learning tasks, where the goal is to minimize the required data. A machine learning method has no adaptability if it requires the same amount of training data in the source and target domains. Prior work suggested that 5000 examples per class is a sufficient rule of thumb for training deep learning architectures from scratch to achieve a minimum acceptable performance [58]. Therefore, an algorithm is considered to have low adaptability if it requires $>$ 2500 examples per class (i.e., many) to optimize a function in a target domain [63]. A 50% reduction in data required for updating a model is non-trivial, and may still constitute a large dataset. An algorithm has high adaptability if it requires $<$ 10 (i.e., a few) examples to update a model, a small dataset in all contexts [62]. It is important to note adaptability is application domain-specific, as the learning tasks are fundamentally different.
IV Non-IID Machine Learning Methods
There are several non-IID machine learning methods that can account for different distribution shift types, including: transfer learning, domain adaptation, test-time training, continual learning, domain generalization, and few-shot learning. Each method corresponds to a specific distribution shift type and the number of distributions considered (see Figure 1). There are two primary distribution shifts: (1) Domain shift, and (2) Label shift. A domain shift indicates a distribution shift of the input features, and label shift indicates a shift of the output values. Label shifts can be characterized by either a new distribution of output values, or a new set of values.
@@ IV Non-IID Machine Learning Methods
Standard IID methods only account for data from a single distribution, which can lead to performance degradation in the presence of distribution shifts [2]. Non-IID methods seek to account for these shifts by intelligently incorporating information from multiple distributions (i.e., domains). Further, these methods either account for exactly two distributions or from more than two distributions, where the boundaries between distributions may be real or artificial. The variability of human behavior is most aptly described by learning tasks with multiple distribution shifts across domains, as humans perform diverse tasks in different ways in a wide range of environments. However, the data can be decomposed or re-organized such that any of the presented techniques can be applied. Data from a single, noisy distribution may be decomposed into multiple distributions to improve generalization to new tasks [64], or data from multiple distributions may be aggregated into a joint distribution to increase data volume and use more complex models [63].
Many of these methods have been successfully applied to task recognition using wearable sensors (e.g., [65, 66, 67, 68]), which shares core commonalities with workload estimation. Thus, these methods provide insight into the capabilities of the underlying machine learning models to account for distribution shifts in relevant application domains. Prior work has primarily focused on evaluating standard IID methods’ ability to generalize to new situations, where many methods fail to generalize to new tasks (e.g., [40, 43, 48, 47]). Relying on standard IID methods ignores the distribution shifts of the physiological signals and workload components as tasks change. No prior work modeled workload estimation for unknown tasks as a non-IID problem. An appropriate non-IID machine learning method for the workload estimation for unknown tasks must be identified, so that the resulting workload estimation algorithm can account for distribution shifts across tasks.
IV-A Domain Adaptation
@@
Domain adaptation considers the setting where training data and test data are characterized by the same features, but are sampled from different distributions.
Definition IV.1
@@
(Domain Adaptation [69]) Given a source domain $D_{S}$ and a learning task $L_{S}$, as well as a target domain $D_{T}$ and learning task $L_{T}$, domain adaptation aims to improve the target predictive function $f_{T}(x)$, where $D_{S}\neq D_{T}$.
This distribution shift type is called a domain shift, or a covariate shift. For example, consider multiple datasets with images of the same objects [70]. Changes in the images’ lighting, background, or orientation represent meaningful difference in the input features and are prototypical examples of a domain shift. Domain adaptation is a non-IID machine learning method that seeks to minimize performance degradation in the presence of domain shift. It is important to note that domain adaptation techniques focus on distribution shifts between one source and one target domain (i.e., two distributions), as shown in Figure 1.
@@
Domain adaptation has also been called transductive transfer learning, as many techniques require the presence of source and target domain data during training in order to perform well [55]. Further, there are two forms of domain adaptation: supervised [70] and unsupervised [61]. Supervised domain requires that labels be available for $L_{T}$, and there are no labels available for $L_{T}$ for unsupervised domain adaptation. There tends to be significantly more data available in source domains than the target domains, and using labeled data from both domains typically results in overfitting to the source distribution [70]. Thus, the majority of prior work has focused on unsupervised domain adaptation.
@@
Domain adaptation has been applied in broad range of HRI application domains, including task recognition (e.g., [71, 72, 65, 73, 74]), workload estimation (e.g., [49, 75, 76]), and emotion recognition (e.g., [77, 78, 79, 80]). The vast majority of these applications use adversarial learning techniques [66, 75, 80], though self-supervised and feature representation learning strategies have also been moderately successful [77].
Adversarial domain adaptation consists of three model types: (1) feature extractor, (2) domain discriminator, and (3) label predictor. Seminal work integrated adversarial learning and domain adaptation as a mini-max game [81]. This work demonstrated that domain adaptation can be achieved by jointly optimizing a feature extraction model to maximize the difference of latent features from source and target domain data, via a domain classifier, and to minimize the error in predicted values for learning tasks, via a label predictor. This joint optimization enables the effective learning of a useful latent feature representation. Prior work demonstrated that replacing the discriminative feature extraction models with generative feature extraction models is also an effective means of domain adaptation (e.g., [82, 83, 84]).
@@
Most domain adaptation techniques exhibit one-to-one portability [84], but multi-target domain adaptation (i.e., one-to-many portability) is an active research area [85]. Domain adaptation techniques primarily rely on deep neural networks [49, 75, 77], consisting of millions of parameters; thus, exhibiting a high model complexity. The adaptability of domain adaptation varies. Some approaches require one example in the target domain [86], while others require thousands [66]. Further, access to source and target data during training is a hard requirement for many domain adaptation techniques, which makes them impractical for real world HRI. One-to-one portability, high model complexity, and low adaptability are undesirable properties; thus, domain adaptation techniques are not viable for workload estimation of unknown tasks.
IV-B Transfer Learning
@@
(Transfer Learning [63]) Given a source domain $D_{S}$, and a learning task $L_{S}$, a target domain $D_{T}$, and learning task $L_{T}$, transfer learning aims to improve the learning of a target predictive function $f_{T}(x)$ in $D_{T}$, using the knowledge in $D_{S}$ and $L_{S}$, where $D_{S}\neq D_{T}$ and $L_{S}\neq L_{T}$.
Transfer learning is characterized a label shift [87], which can refer to a shift in the relative distribution of output values (e.g., disease prevalence in a rural hospital vs. urban hospital), or a complete change in the labels available. Transfer learning techniques are capable of transferring knowledge from a model trained to identify everyday objects (e.g., cats, planes) in images to a model performing medical image analysis (e.g., ultrasounds, x-rays) [88]. Similar to domain adaptation, transfer learning has been widely applied to HRI (e.g., human task recognition [89, 90], emotion recognition [91, 92, 93, 94]). There are many transfer learning techniques, which can be categorized based on how knowledge is transferred: instance-based (e.g., [95, 96, 97]), feature-based (e.g., [98, 99, 100]), and parameter-based (e.g., [101, 87, 68, 102]). Instanced and feature-based techniques have had moderate levels of success, but are seldom used in modern applications when compared to parameter-based techniques [103].
@@
Parameter-based methods have two phases: pre-training and fine-tuning. Pre-training trains a machine learning model on data from the source domain, $D_{S}$, to solve the source learning task $L_{S}$. Fine-tuning trains the same model on data from the target domain, $D_{T}$, but only updates a subset of the parameters. This two-step training procedure enables the network to learn a useful latent representation during pre-training, and a task-specific transformation of that representation during fine-tuning. This approach has been successful in computer vision [87], robotics [104], and task recognition (e.g., [105, 106, 107, 108]).
@@
Many of these methods use deep learning architectures to transfer knowledge across source/target domain pairs, requiring thousands of examples to be successful. Further, there is evidence to suggest that the model complexity is inversely related to both portability and adaptability [63] [59]. Zhuang et al. evaluated a range of deep learning architectures on several tasks, all of which exhibited one-to-one portability, high model complexity, and low adaptability [63]. Similar findings exist in transfer learning for human activity recognition [105]. Similar techniques were also developed to solve a wide range of natural language problems based on only a few examples [59]. This high adaptability is in direct contrast to the previously discussed transfer learning techniques, where the primary difference is the model complexity. A transformer deep learning architecture was trained to predict the next word in a sentence [59]. This network was the foundation of several downstream networks that performed tasks, including machine translation, question answering, and arithmetic word problem-solving. The architecture consisted of 175 billion parameters, requiring significantly more data to train parameter-based techniques that rely on smaller networks (i.e., tens of billions of data points). These results suggest that a higher model complexity can directly improve portability and adaptability.
@@
A core limitation of workload estimation is data volume, as gathering large volumes of wearable sensor data is logistically impractical; thus, model complexity is constrained. It is also desirable that workload estimation algorithms exhibit many-to-many portability, as models may need information across multiple individuals and tasks. Thus, transfer learning is not a viable option for workload estimation of unknown tasks.
IV-C Test-time Adaptation

### v2

@@ 3.2 Evaluation Criteria
Portability refers to the number of distributions used to train and test a model (Sun et al., 2015). Learning tasks may consist of multiple source domains (e.g., urban vs wildland fire fighting). A straightforward means of combining information from multiple domains is to construct a single joint distribution; however, data in some domains may be of higher quality than others. Weighting high-quality data over low-quality data is appropriate to maximize performance (Sun et al., 2015). Additionally, the separation of source domains may be artificial. Prior work established that modeling domains separately can be beneficial for learning inductive biases that improve performance on unknown learning tasks (Baxter, 2000). Portability can have four values: one-to-one, one-to-many, many-to-one, and many-to-many. One-to-one refers to an algorithm that transfers knowledge between one source and one target domain. One-to-many transfers knowledge from one source domain to many target domains, and many-to-one transfers knowledge from multiple source domains to a single target domain. Many-to-many knowledge transfer occurs between multiple sources to multiple target domains simultaneously.
A many-to-many portability is desirable for workload estimation for unknown tasks. Machine learning-based workload estimation methods are primarily developed using data collected in experimental conditions, producing diverse datasets with low data volumes. Further, these experimental conditions may unintentionally introduce bias into the dataset (Bhagat Smith et al., 2024). Developing workload estimation methods must therefore be able to accommodate distribution shifts for multiple, small, noisy datasets (i.e., many-to-many portability). Many-to-one portability is only desirable if a method needs to estimate workload for a single new task, which is unlikely in real-world application domains. One-to-one portability is not sufficient outside of contrived circumstances, as effectively capturing the nuanced differences between known and unknown tasks using two distributions requires large data volumes.
@@
Figure 1: A high level overview of non-IID machine learning methods. Adapted from Figure 19.11 of (Murphy, 2023, chap. 19).
@@
Model Complexity refers to the number of parameters an underlying machine learning model needs for the algorithm to be successful (Goodfellow et al., 2016). Machine learning models with more parameters require higher volumes of training data. Additionally, training machine learning models in human-robot interaction (HRI) application domains requires this data to be ecologically valid, introducing additional complexity to model development. Early statisticians built simple linear models using hundreds of examples (Goodfellow et al., 2016), whereas modern deep learning architectures may require tens of billions of examples (Brown et al., 2020). Model complexity shares a relationship with both the underlying machine learning model type and problem complexity, as complex models are often required to solve complex problems. Workload estimation and many HRI problems are typically conducted in low-data regimes, as they require human subject evaluations for data collection; thus, smaller models are preferred. An example of low model complexity is a random forest model that contains a few hundred decision trees and typically contain on the order of thousands of parameters (Breiman, 2001). An example of medium model complexity is shallow neural networks (i.e., fully connected neural networks) that typically contain hundreds of thousands of parameters (Fortune et al., 2020), but may contain on the order of millions of parameters. Two examples of high model complexity are deep learning and generative AI architectures (e.g., convolutional networks, long short-term memory networks (LSTM), transformers). These architectures can contain anywhere from tens of millions (Wilson and Cook, 2020) to over 100 billion parameters (Brown et al., 2020).
A machine learning-based workload estimation method must be capable of running onboard a robot to be useful in uncertain, dynamic environments. The robot is not guaranteed to have a remote connection to a powerful computer or access to the cloud. Modern edge compute solutions (e.g., NVIDIA Jetson Orin (NVIDIA, 2025)) have the capability of executing many deep learning architectures in real-time, but it is desirable to minimize the model complexity in order to reduce the requisite computational resources and eliminate reliance on cloud resources. Further, workload estimation methods will be incorporated into a larger human-robot teaming architecture that may rely on other machine learning models. Reducing the model complexity of the workload estimation method frees up resources required by other components. Minimizing model complexity reduces the power consumed by the machine learning methods allow for more efficient power consumption, a critical constraint in uncertain, dynamic environments. Smaller models also require fewer computational resources that may be necessary for other learning tasks (e.g., task recognition, decision making).
@@
The last metric, Adaptability, refers to the amount of data required to update a machine learning model for new learning tasks (Wang et al., 2020b). Many non-IID methods seek to account for distribution shifts by updating existing machine learning models’ parameters. Adaptability is the number of test data points necessary for the updated model to make accurate predictions for new learning tasks, where the goal is to minimize the required data. A machine learning method has no adaptability if it requires the same amount of training data in the source and target domains. Prior work suggested that 5000 examples per class is a sufficient rule of thumb for training deep learning architectures from scratch to achieve a minimum acceptable performance (Goodfellow et al., 2016). A 50% reduction in data required for updating a model is non-trivial, and may still constitute a large dataset. Therefore, an algorithm is considered to have low adaptability if it requires $>$ 2500 examples per class (i.e., many) to optimize a function in a target domain (Zhuang et al., 2020). An algorithm has high adaptability if it requires $<$ 10 (i.e., a few) examples to update a model, a small dataset in all contexts (Wang et al., 2020b).
Adaptability is application-domain specific, as learning tasks are fundamentally different. Real-world human-robot teams will encounter unknown tasks in the field and must be able to estimate the human’s workload as quickly as possible. Achieving a high adaptability is ideal, requiring a workload estimation method to collect large amounts of data in dynamic environments is impractical.
4 Non-IID Machine Learning Methods
Non-IID machine learning methods seek to account for distribution shifts by intelligently incorporating information from multiple distributions (i.e., domains) and fall into six categories: domain adaptation, transfer learning, test-time training, continual learning, domain generalization, and few-shot learning (see Figure 1). These categories are primarily differentiated by their data handling procedure, which can also be viewed as how each method formulates the distribution shift problem (Murphy, 2023, chap. 19). Analyzing each category’s data handling procedures is central to understanding a method’s practical utility for estimating workload associated with unknown tasks, as it provides insight into how these data handling procedures can take advantage of the unique characteristics of HRI datasets.
Datasets used to train machine learning-based workload estimation methods are heavily impacted by the variability inherent in the physiological signals, noise within wearable sensor data, and range of potential human behaviors (Das Chakladar and Roy, 2024). These data characteristics present challenges unique to HRI application domains and differentiate them from other learning tasks (e.g., object detection, speech recognition). Additionally, most machine learning-based workload estimation methods rely on datasets collected in controlled experimental conditions.
Methods within each category implement similar data handling procedures, where each category varies based upon the distribution shift type and number of distributions considered. There are two primary distribution shift types. A domain shift indicates a distribution shift of the input features (e.g., heart-rate variability, pupil diameter), and label shift (e.g., workload value) indicates a shift of the output values. Domain shifts can be characterized by changes to the range of values in various physiological signals or changes to the relative importance for any given physiological signal for estimating workload. Label shifts can be characterized by either a new distribution of output values, or a new set of values. For example, consider a training dataset where the cognitive workload values range from 0 to 15, with a mean value of 10 and a median value of 5. A testing dataset with a label shift may have cognitive workload values ranging from 5 to 20, or may have a mean and median value of 14. The degree to which these distribution shifts impact performance varies based upon the shift’s magnitude and dataset size.
@@ 4 Non-IID Machine Learning Methods
Non-IID machine learning methods either account for exactly two distributions or more than two distributions, where the boundaries between distributions may be real or artificial. Human behavior is most aptly described by learning tasks with multiple distribution shifts across domains, as humans perform diverse tasks in different ways in a wide range of environments. However, the data can be decomposed or re-organized such that any of the presented techniques can be applied. Data from a single, noisy distribution may be decomposed into multiple distributions to improve generalization to new tasks (Vinyals et al., 2016), or data from multiple distributions may be aggregated into a joint distribution to increase data volume and use more complex models (Zhuang et al., 2020).
An analysis of non-IID machine learning methods that considers the unique characteristics of HRI datasets and how they interact with each method’s data handling procedure is presented. Relevant methods within each category are briefly discussed to assess their merit for estimating human workload for unknown tasks using the developed criteria (i.e., portability, model complexity, and adaptability).
4.1 Domain Adaptation
@@
Domain adaptation considers scenarios where training and testing data contain the same features (e.g., physiological signals), but are sampled from different distributions.
Definition:.
@@
(Domain Adaptation) Given a source domain $D_{S}$ and a learning task $L_{S}$, as well as a target domain $D_{T}$ and learning task $L_{T}$, domain adaptation aims to improve the target predictive function $f_{T}(x)$, where $D_{S}\neq D_{T}$ (Singhal et al., 2023).
This distribution shift type is called a domain shift, or a covariate shift. An intuitive example is a machine learning-based workload estimation method trained on data collected in an climate controlled environment (i.e., source domain) and tested on data in an outdoor environment (i.e., target domain). Extreme temperature differences will have an impact on an individual’s workload, but also alter their physiological responses to nominal workload conditions (Hancock and Vasmatzidis, 2003). Domain adaptation is a non-IID machine learning method that seeks to minimize performance degradation in the presence of domain shift. It is important to note that domain adaptation techniques focus on distribution shifts between one source and one target domain (i.e., two distributions), as shown in Figure 1.
@@
Domain adaptation has also been called transductive transfer learning, as many techniques require the presence of source and target domain data during training to perform well (Pan and Yang, 2009). Further, there are two forms of domain adaptation: supervised (Wang and Deng, 2018) and unsupervised (Wilson and Cook, 2020). Supervised domain adaptation requires that labels be available for $L_{T}$, while there are no labels available for $L_{T}$ for unsupervised domain adaptation. There tends to be significantly more data available in source domains, and using labeled data from both domains typically results in overfitting to the source distribution (Wang and Deng, 2018). Thus, the majority of prior work has focused on unsupervised domain adaptation.
@@
Domain adaptation has been applied in broad range of HRI application domains, including task recognition (e.g., Alajaji et al., 2023; An et al., 2021; Faridee et al., 2022; Zakia and Menon, 2021; Li et al., 2022b), workload estimation (e.g., Zhou et al., 2022, 2023b; Albuquerque et al., 2022), and emotion recognition (e.g., Li et al., 2022c; Latif et al., 2022; Guo et al., 2023; He et al., 2022). The vast majority of these applications use adversarial learning techniques (Li et al., 2021c; Zhou et al., 2023b; He et al., 2022), though self-supervised and feature representation learning strategies have also been moderately successful (Li et al., 2022c).
@@
Most domain adaptation techniques exhibit one-to-one portability (Balaji et al., 2020; Liu and Tuzel, 2016; Li et al., 2020), but multi-target domain adaptation (i.e., one-to-many portability) is an active research area (Gholami et al., 2020). Domain adaptation techniques primarily rely on deep neural networks (Zhou et al., 2022, 2023b; Li et al., 2022c) consisting of millions of parameters; thus, exhibiting a high model complexity. Domain adaptation techniques exhibit varying adaptability, where some approaches require one example in the target domain (Yue et al., 2021), while others require thousands (Li et al., 2021c). Further, access to source and target data during training is a core requirement for all domain adaptation techniques, making them impractical for real-world HRI. One-to-one portability, high model complexity, and low adaptability are undesirable properties; thus, domain adaptation techniques are not viable for workload estimation of unknown tasks.
4.2 Transfer Learning
@@
(Transfer Learning) Given a source domain $D_{S}$, and a learning task $L_{S}$, a target domain $D_{T}$, and learning task $L_{T}$, transfer learning aims to improve the learning of a target predictive function $f_{T}(x)$ in $D_{T}$, using the knowledge in $D_{S}$ and $L_{S}$, where $D_{S}\neq D_{T}$ and $L_{S}\neq L_{T}$ (Zhuang et al., 2020).
Transfer learning is characterized by a label shift (Huh et al., 2016), either in the relative distribution of output values or a complete change in the available labels. Consider the field of affective computing’s use of physiological signals (Egger et al., 2019). Transfer learning techniques attempt to transfer knowledge from a model trained to recognize positive emotions (e.g., awe) to a model attempting to recognize negative emotions (e.g., anger, disgust). These techniques are widely applied within HRI application domains (e.g., human task recognition (Ding et al., 2018; Li et al., 2021b), emotion recognition (Li et al., 2019a; Quan et al., 2023; Ma et al., 2023; Nguyen et al., 2023)).
There are many transfer learning techniques that vary based on how knowledge is transferred: instance-based (e.g., Chen et al., 2022c; Wang et al., 2019a, b), feature-based (e.g., Ren et al., 2019; Stüber et al., 2018; Wu et al., 2021), and parameter-based (e.g., Fu et al., 2021; Huh et al., 2016; Soleimani and Nazerfard, 2021; Link et al., 2022). Instanced and feature-based techniques have had moderate levels of success, but are seldom used in modern applications when compared to parameter-based techniques (Weiss et al., 2016).
@@
Parameter-based methods have two phases: pre-training and fine-tuning. Pre-training optimizes a machine learning model on data from the source domain, $D_{S}$, to solve the source learning task $L_{S}$. Fine-tuning trains the same model on data from the target domain, $D_{T}$, but only updates a subset of the parameters. This two-step training procedure enables the network to learn a useful latent representation during pre-training, and a task-specific transformation of that representation during fine-tuning. This approach has been successful in computer vision (Huh et al., 2016), robotics (Lee et al., 2021), and task recognition (e.g., An et al., 2023; Ray et al., 2023; Pavliuk et al., 2023).
@@
Many of these methods use deep learning architectures to transfer knowledge across source/target domain pairs, requiring thousands of examples to be successful. Further, there is evidence to suggest that the model complexity is inversely related to both portability and adaptability (Zhuang et al., 2020; Brown et al., 2020). Zhuang et al. evaluated a range of deep learning architectures on several tasks, all of which exhibited one-to-one portability, high model complexity, and low adaptability (Zhuang et al., 2020). Similar techniques were developed to solve a wide range of natural language problems based on only a few examples (Brown et al., 2020). This high adaptability is in direct contrast to the other transfer learning techniques, where the primary difference is the model complexity. A transformer deep learning architecture was trained to predict the next word in a sentence (Brown et al., 2020). This network was the foundation of several downstream networks that performed tasks, including machine translation, question answering, and arithmetic word problem-solving. The architecture consisted of 175 billion parameters, requiring significantly more data to train than the parameter-based techniques that use smaller networks (i.e., tens of billions of data points). These results suggest that a higher model complexity can directly improve portability and adaptability.
@@
A core limitation of workload estimation is data volume, as gathering large volumes of wearable sensor data is logistically impractical; thus, model complexity is constrained. It is also desirable that workload estimation algorithms exhibit many-to-many portability, as models may need information across multiple individuals and tasks. Thus, transfer learning is not a viable option.
4.3 Test-time Adaptation